Circular No. 8271

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ls NEAR PGC 11402 Further to IAUC 8266, R. J. Foley and A. V. Filippenko, University of

California at Berkeley (UCB), report the LOSS discovery, on KAIT images taken on 2003 Dec. 22.3 (mag ∼ 16.5) and 28.2 UT (mag ∼ 17.0), of an apparent supernova located at α = 3^{h}01^{m}00^{s}.39, δ = −10^{o}53′04′′.5 (equinox 2000.0; revision to position given on CBET 58), which is 170′′ west and 55′′ north of the nucleus of PGC 11402. A KAIT image taken on Nov. 25.3 showed nothing at this position (limiting mag ∼ 19.0). Foley, F. J. D. Serduke (Lawrence Livermore National Laboratory), R. Chornock (UCB), and Filippenko report that inspection of CCD spectra (range 330–1000 nm), obtained on 2004 Jan. 17 with the Shane 3-m telescope at Lick Observatory, reveals that the new object is of type Ia, ∼ 1.5 months past maximum brightness. Its redshift, estimated by comparing the spectrum with that of other type-Ia supernovae, is ∼ 13000 km/s — much greater than that of PGC 11402 (9276 km/s, from NED), so it must be in a faint background galaxy unrelated to PGC 11402.

SUPERNOVAE 2004D, 2004E, AND 2004F B. Swift, J. Burket, H. Pugh, R. J. Foley, A. V. Filippenko, and W. Li

report the LOSS discovery of three additional supernovae on KAIT images.

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004D Jan. 4.5 11^{h}56^{m}27.^{s}82 +39^{o}44′16′′.7 19.1 4′′.7 W, 14′′.1 S
2004E Jan. 15.5 13 16 51.40 +31 35 11.8 16.1 3′′.2 E, 20′′.4 N
2004F Jan. 16.2 3 17 53.80 − 7 17 43.0 17.8 6′′.0 E, 9′′.2 N

Additional KAIT magnitudes: SN 2004D in UGC 6916, 2003 May 22.3 UT, [19.5; 2004 Jan. 10.5, 19.0. SN 2004E in PGC 46239, 2002 Dec. 25.1, [19.5; 2004 Jan. 16.5, 16.1. SN 2004F in NGC 1285, 2003 Nov. 21.2, [18.5; 2004 Jan. 17.2, 17.9. Foley, Serduke, Chornock, and Filippenko add that inspection of CCD spectra, obtained as above on Jan. 17, reveals that SN 2004D is of type II, several months past maximum brightness; relatively narrow Hα emission is dominant, and the Ca II near-infrared triplet emission is also visible, but [O I] and [Ca II] are weak or nonexistent, so the object has not yet reached the nebular phase. SN 2004E is of type Ia, with a spectral-feature age (Riess et al. 1997, A.J. 114, 722) of 2 ± 2 days past maximum brightness.

2004 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green